One ordered pair of rational numbers (x,y) satisfies √((21/4) + 3√3) = x + √y. What is the value of x + y?
Answer by ewatrrr(24785)


Hi
One ordered pair of rational numbers (x,y) satisfies:
√(2 1/4) + 3√3) = x + √y.
√(9/4) + 3√3 = x + √y.
3/2 + 3√3) = x + √y.
x = 3/2
√y = 3√3 0r squaring both sides
y = 9*3 = 27
x + y = 3/2 + 27 = 28 1/2
Wish You the Best in your Studies.